OpenStudy (yisselluna):

Find the product. x 5 · x 4 · x 3

OpenStudy (sooobored):

\[x^a*x^b*x^c=x^{a+b+c}\] \[ax*bx*cx = a*b*c*x^3\]

OpenStudy (yisselluna):

so its x^3??

OpenStudy (sooobored):

5+4+3 is not equal to 3